Let and be two functions defined by
F(x)=x^2+x-12
g(x)=x^2-7x-8.
Find all values of x that are NOT in the domain of the function f/g.?

Answer by nerdybill(7384)

Domain specifies the "values" 'x' can take.
.
F(x)=x^2+x-12
g(x)=x^2-7x-8.
Find all values of x that are NOT in the domain of the function f/g.?
.
f/g is therefore:
(x^2+x-12)/(x^2-7x-8)
The ONLY values x CAN'T take are values that make the denominator go to zero.
Look for those by setting the denominator to zero and solving for x:
x^2-7x-8 = 0
(x-8)(x+1) = 0
x = {-1, 8}
So, the ONLY two values 'x' can't take is -1 and 8.
Therefore, the domain is:
(-oo, -1) U (-1, 8) U (8,+oo)
OR, you could say:
x: x all real numbers except -1 and 8